# From harmonic to Newman-Unti coordinates at the second post-Minkowskian order

### v1

@@
Abstract: In this paper, we present the complete transformations of a generic metric from harmonic to Newman-Unti coordinates up to the second post-Minkowskian order $(G^{2})$. This allows us to determine the asymptotic shear, the Bondi mass aspect, and the angular-momentum aspect at both orders.
@@ 1 Introduction
In this paper, we adopt a slightly different strategy to derive the concrete coordinate transformations. In [24, 25], the transformation between the two coordinate systems was constructed by expressing the NU coordinates in terms of the harmonic coordinates, which manifests the relationship between the two coordinate systems. However, to extract the asymptotic data of the metric in the NU gauge, one must then invert these relations, since the original metric is given in harmonic coordinates. In contrast, we assume in this work that the harmonic coordinates are expressed in terms of the NU coordinates. Furthermore, we assume that the perturbative metric in harmonic coordinates admits an expansion in inverse powers of the radial coordinate at large distances. Under these assumptions, we explicitly compute the coordinate transformations that map a generic perturbative metric into the NU gauge at the first and second PM orders. The asymptotic forms of the perturbative metric in the NU gauge is obtained directly from the change of the coordinates. In particular, we identify the asymptotic data, such as the asymptotic shear, Bondi mass aspect, and angular momentum aspect, from the metric in NU gauge. Our results can be generally applied to any solutions given in the harmonic gauge.
@@ 2 Algorithm of the coordinate transformation and the NU metric at linear order
Note that those conditions are weaker than smoothness of the boundary metric at null infinity, see, e.g., [38]. Hence, there will be logarithmic terms in the metric components in the NU gauge. We will not justify the physical consequence of the above fall-off conditions. Nevertheless, a large class of matter fields preserving asymptotically flat conditions must satisfy those conditions, such as Maxwell fields, scalar fields, or massive moving particles. The constraints from the conditions in (8) will be applied to simplify the metric expressions in the NU gauge. [^2] The algorithm developed in this work can, in principle, be applied without imposing the conditions in (8). However, the resulting metric at order $G^{2}$ in the NU gauge is very complicated with logarithmic terms appearing in all metric components. For simplicity, we apply the conditions in (8) which still encompass most physically interesting systems. The linearized Einstein equation yields that
@@
where $U_{10}$, $R_{10}$, and $X_{10}^{A}$ are integration constants with respect to the radial variable $r$. Above, we only present the orders that are relevant to asymptotic data in the NU framework in the $1/r$ expansion. The fall-off conditions of the NU gauge are
@@
and [^3] In the standard Bondi or NU gauge, the angular momentum aspect is not the full $g_{uA}$ component at order $1/r$, see, e.g., [21, 23]. However, the difference is only a total derivative term which does not affect the definition of angular momentum. Moreover, we only introduce the constraints from the fall-off conditions of the stress-energy tensor. The perturbative metric does not necessarily satisfy the Einstein equation. So the metric form in the NU gauge may not fully align with the results in [21, 23].
@@ 3 Metric in NU gauge at quadratic order
In this section, we will derive the metric in NU gauge at order $G^{2}$. Before computing the coordinate transformations, we first revisit the fall-off conditions for the stress-energy tensor imposed in (8). At order $G^{2}$, the stress-energy tensor contains two contributions: the matter stress–energy tensor and the effective stress–energy tensor induced by the order $G$ perturbative metric. We find that the effective stress–energy tensor constructed from the general asymptotic form of the perturbative metric in (6) does not fully obey the fall-off conditions in (8). Therefore, for a consistent computation at order $G^{2}$, we must impose stronger fall-off conditions at ${\cal O}(G)$, namely, [^4] The algorithm developed in this work can be definitely applied to the computation at ${\cal O}(G^{2})$ without imposing these extra conditions. Correspondingly, the relations in (9) are not valid at ${\cal O}(G^{2})$. However, imposing those conditions can lead to a consistent computation for both PM orders, which could manifest some generic features at different PM orders.
@@
$\partial_{r}U_{2}=$ $\frac{1}{2}{h^{(f)}_{2}}_{r_{f}r_{f}}-\frac{1}{2}(\partial_{r}U_{1})^{2}-\partial_{r}U_{1}\partial_{r}R_{1}+\frac{1}{2}r^{2}\bar{\gamma}_{AB}\partial_{r}X_{1}^{A}\partial_{r}X_{1}^{B}+{h^{(f)}_{1}}_{r_{f}u_{f}}\partial_{r}U_{1}+{h^{(f)}_{1}}_{r_{f}r_{f}}\partial_{r}R_{1}$
$+{h^{(f)}_{1}}_{r_{f}A}\partial_{r}X^{A}_{1}+\frac{1}{2}U_{1}\partial_{u}{h^{(f)}_{1}}_{r_{f}r_{f}}+\frac{1}{2}R_{1}\partial_{r}{h^{(f)}_{1}}_{r_{f}r_{f}}+\frac{1}{2}X^{A}_{1}D_{A}{h^{(f)}_{1}}_{r_{f}r_{f}},$ (35)
$\partial_{r}R_{2}=$ ${h^{(f)}_{2}}_{u_{f}r_{f}}-\partial_{r}U_{1}\partial_{u}U_{1}-\partial_{r}U_{1}\partial_{u}R_{1}-\partial_{r}R_{1}\partial_{u}U_{1}+r^{2}\bar{\gamma}_{AB}\partial_{r}X_{1}^{A}\partial_{u}X_{1}^{B}+{h^{(f)}_{1}}_{r_{f}u_{f}}\partial_{u}U_{1}$
$+{h^{(f)}_{1}}_{r_{f}r_{f}}\partial_{u}R_{1}+{h^{(f)}_{1}}_{r_{f}A}\partial_{u}X_{1}^{A}+{h^{(f)}_{1}}_{u_{f}u_{f}}\partial_{r}U_{1}+{h^{(f)}_{1}}_{u_{f}r_{f}}\partial_{r}R_{1}+{h^{(f)}_{1}}_{u_{f}A}\partial_{r}X_{1}^{A}$
$+U_{1}\partial_{u}{h^{(f)}_{1}}_{u_{f}r_{f}}+R_{1}\partial_{r}{h^{(f)}_{1}}_{u_{f}r_{f}}+X_{1}^{A}D_{A}{h^{(f)}_{1}}_{u_{f}r_{f}}-\partial_{r}U_{2}-\partial_{u}U_{2},$ (36)
@@
$+U_{1}\partial_{u}{h^{(f)}_{1}}_{{r_{f}B}}+R_{1}\partial_{r}{h^{(f)}_{1}}_{{r_{f}B}}+X_{1}^{C}D_{C}{h^{(f)}_{1}}_{{r_{f}B}}+{h^{(f)}_{1}}_{Bu_{f}}\partial_{r}U_{1}+{h^{(f)}_{1}}_{Br_{f}}\partial_{r}R_{1}$
$+{h^{(f)}_{1}}_{BC}\partial_{r}X_{1}^{C}-\partial_{r}U_{1}D_{B}U_{1}-\partial_{r}U_{1}D_{B}R_{1}-\partial_{r}R_{1}D_{B}U_{1}-D_{B}U_{2}+2rR_{1}\partial_{r}X_{1B}\bigg)$
$-\bar{\gamma}^{AB}\partial_{r}X_{1C}\partial_{B}X_{1}^{C}+X_{1}^{C}\partial_{C}\bar{\gamma}^{AB}\partial_{r}X_{1B}.$ (37)
@@
$+\frac{1}{2r^{2}}\bigg[{{h_{21}^{(f)}}_{r_{f}}}^{A}+{h_{11}^{(f)}}_{u_{f}r_{f}}D^{A}\beta_{1}-{{h_{1m}^{(f)}}_{B}}^{A}\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)-2R_{10}\left({{h_{10}^{(f)}}_{r_{f}}}^{A}-D^{A}\beta_{1}\right)$
$\qquad-D^{A}U_{21}+\beta_{1}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{A}+\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)D_{B}{{h_{10}^{(f)}}_{r_{f}}}^{A}+{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{A}\left({{h_{10}^{(f)}}_{r_{f}B}}-D_{B}\beta_{1}\right)$
@@
$\qquad+\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{A}\left({{h_{11}^{(f)}}_{r_{f}B}}+\frac{1}{2}D_{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)+\frac{1}{2}\left({{h_{11}^{(f)}}_{r_{f}}}^{B}+\frac{1}{2}D^{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)D_{B}{{h_{10}^{(f)}}_{r_{f}}}^{A}$
$\qquad+{{h_{11}^{(f)}}_{r_{f}B}}D^{A}\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)+\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)D_{B}{{h_{11}^{(f)}}_{r_{f}}}^{A}$
@@
$\qquad{C_{2}}_{AB}={h_{2m}^{(f)}}_{AB}+2D_{(A}{h_{20}^{(f)}}_{B)r_{f}}-D_{A}D_{B}{h_{21}^{(f)}}_{r_{f}r_{f}}-2D_{A}D_{B}\beta_{2}+\bar{\gamma}_{AB}D^{2}\beta_{2}$ (47)
$\qquad\qquad\qquad-\frac{1}{2}\bar{\gamma}_{AB}\bar{\gamma}^{CD}{h_{2m}^{(f)}}_{CD}-\bar{\gamma}_{AB}D^{C}{h_{20}^{(f)}}_{r_{f}C}+\frac{1}{2}\bar{\gamma}_{AB}D^{2}{h_{21}^{(f)}}_{r_{f}r_{f}},$
@@
$+\frac{1}{3}D^{B}\beta_{1}\partial_{u}\left({h_{10}^{(f)}}_{AB}+D_{(A}{h_{10}^{(f)}}_{B)r_{f}}+\frac{1}{2}D_{A}D_{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)+D^{B}\beta_{1}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)$
$+\frac{1}{2}D^{B}\beta_{1}D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)+\frac{1}{2}D^{B}\beta_{1}D^{C}\left(D_{A}{h_{1m}^{(f)}}_{BC}-D_{B}{h_{1m}^{(f)}}_{AC}\right)$
@@
$-\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{C}\left(D_{A}{h_{1m}^{(f)}}_{BC}-D_{B}{h_{1m}^{(f)}}_{AC}\right)-\frac{1}{3}{{h_{10}^{(f)}}_{r_{f}}}^{B}\left(\partial_{u}{h_{10}^{(f)}}_{AB}+\partial_{u}D_{(A}{h_{10}^{(f)}}_{B)r_{f}}\right)$
$-\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)-{{h_{10}^{(f)}}_{r_{f}}}^{B}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)$
$+\frac{1}{6}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{B}\left({h_{1m}^{(f)}}_{AB}+D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right).$ (49)
@@
$D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)\\ =2D^{C}\left(D_{A}D_{(C}{h_{10}^{(f)}}_{B)r_{f}}-D_{B}D_{(C}{h_{10}^{(f)}}_{A)r_{f}}\right)+2\left(D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right).$ (50)
@@
${N_{2}}_{A}={h_{21}}_{uA}=$ ${h_{21}^{(f)}}_{u_{f}A}+D_{A}{h_{22}^{(f)}}_{u_{f}r_{f}}+\frac{1}{3}\partial_{u}{h_{22}^{(f)}}_{r_{f}A}+\frac{1}{3}\partial_{u}D_{A}{h_{23}^{(f)}}_{r_{f}r_{f}}$
@@
$+\frac{1}{3}(D^{B}\beta_{1}-{{h_{10}^{(f)}}_{r_{f}}}^{B})\partial_{u}\left({h_{10}^{(f)}}_{AB}+D_{(A}{h_{10}^{(f)}}_{B)r_{f}}\right)$
@@
$+\frac{1}{6}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{B}\left({h_{1m}^{(f)}}_{AB}+D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right),$ (52)

### v2

@@
Abstract: In this paper, we present the complete transformations of a generic metric from (generalized) harmonic to Newman-Unti coordinates up to the second post-Minkowskian order $(G^{2})$. This allows us to determine the asymptotic shear, the Bondi mass aspect, and the angular-momentum aspect at both orders.
@@ 1 Introduction
In this paper, we adopt a slightly different strategy to derive the concrete coordinate transformations. In [24, 25], the transformation between the two coordinate systems was constructed by expressing the NU coordinates in terms of the harmonic coordinates, which manifests the relationship between the two coordinate systems. However, to extract the asymptotic data of the metric in the NU gauge, one must then invert these relations, since the original metric is given in harmonic coordinates. In contrast, we assume in this work that the harmonic coordinates are expressed in terms of the NU coordinates. Furthermore, we assume that the perturbative metric in harmonic coordinates admits an expansion in inverse powers of the radial coordinate at large distances $r$. In standard harmonic coordinates, logarithmic terms arise naturally in the solution of linearized Einstein equation satisfying the harmonic gauge condition. For example, such terms appear at the third post-Newtonian order in the two-body problem [37]. In contrast, we are working with generalized harmonic coordinates under the assumption of a large-distance expansion in integer powers of $1/r$. Nevertheless, logarithmic terms can always be eliminated to arbitrarily high order in $1/r$ expansion by suitably generalizing the harmonic gauge conditions [38]. Under these assumptions, we explicitly compute the coordinate transformations that map a generic perturbative metric into the NU gauge at the first and second PM orders. The asymptotic forms of the perturbative metric in the NU gauge is obtained directly from the change of the coordinates. In particular, we identify the asymptotic data, such as the asymptotic shear, Bondi mass aspect, and angular momentum aspect, from the metric in NU gauge. Our results can be generally applied to any solutions given in the harmonic gauge.
@@ 2 Algorithm of the coordinate transformation and the NU metric at linear order
It is important to clarify that we are not working with standard harmonic coordinates, since we do not impose the corresponding harmonic gauge conditions on the perturbative metric. Rather, the coordinates are harmonic only with respect to the flat background. The constraints on the perturbative metric are that it admits a large-distance expansion in integer powers of $1/r_{f}$. Such conditions can always be achieved within the framework of generalized harmonic coordinates for solutions of linearized Einstein equation [38].
@@
Note that those conditions are weaker than smoothness of the boundary metric at null infinity, see, e.g., [39]. Hence, there will be logarithmic terms in the metric components in the NU gauge. In the PM expansion, the stress-energy tensor at higher orders in $G$ (starting at $G^{2}$) consists of two contributions: one from the matter fields and another from the lower-order perturbative metric. The latter part is referred to as effective stress-energy tensor. The fall-off conditions in (8) are intended to hold generally for the full stress-energy tensor at all orders in $G$. This requirement can impose additional constraints on the asymptotic behavior of the perturbative metric, ensuring that the effective stress-energy tensor satisfies the fall-off conditions in (8). These constraints will be discussed explicitly at order $G^{2}$ at the beginning of the next section. In particular, the effective stress-energy tensor at order $G^{2}$, which guarantees the smoothness of the boundary metric at null infinity, imposes strong restrictions on the order $G$ metric, which would rule out many physically interesting solutions. Thus we adopt a weaker fall-off conditions for the stress-energy tensor than the smoothness conditions. The constraints from the conditions in (8) will be applied to eventually simplify the metric expressions in the NU gauge. [^2] The algorithm developed in this work can, in principle, be applied without imposing the conditions in (8). However, the resulting metric at order $G^{2}$ in the NU gauge is very complicated with logarithmic terms appearing in all metric components. For simplicity, we apply the conditions in (8) which still encompass most physically interesting systems. Recalling that the fall-off conditions for the stress-energy tensor is imposed generically, including the effective stress-energy tensor from the metric at lower orders, the linearized Einstein equations at all PM orders should be consistent with those fall-off conditions, which yields that
@@
where $U_{10}$, $R_{10}$, and $X_{10}^{A}$ are integration constants with respect to the radial variable $r$. Above, we only present the orders that are relevant to asymptotic data in the NU framework in the $1/r$ expansion. The fall-off conditions of the NU gauge are [8, 9]
@@
and [^3] In the standard Bondi or NU gauge, the angular momentum aspect is not the full $g_{uA}$ component at order $1/r$, see, e.g., [21, 23]. However, the difference is only a total derivative term which does not affect the definition of angular momentum. In this work, we only introduce some fall-off conditions for the stress-energy tensor without specifying its precise formulas. Hence, the perturbative metric is only a solution to the linearized Einstein equation to the orders with respect to the fall-off conditions of the stress-energy tensor in $\frac{1}{r}$ expansion. The exact expressions of the obtained metric components in NU gauge may not fully align with the results in [21, 23] where the metric is a solution of Einstein equation to all orders.
@@ 3 Metric in NU gauge at quadratic order
In this section, we will derive the metric in NU gauge at order $G^{2}$. Before computing the coordinate transformations, we first revisit the fall-off conditions for the stress-energy tensor imposed in (8). At order $G^{2}$, the stress-energy tensor contains two types of contribution, namely, the matter stress–energy tensor and the effective stress–energy tensor induced by the order $G$ perturbative metric. We find that the effective stress–energy tensor constructed from the general asymptotic form of the perturbative metric in (6) does not fully obey the fall-off conditions in (8). Therefore, for a consistent computation at order $G^{2}$, we must impose stronger fall-off conditions at ${\cal O}(G)$, namely, [^4] The algorithm developed in this work can be definitely applied to the computation at ${\cal O}(G^{2})$ without imposing these extra conditions. Correspondingly, the relations in (9) are not valid at ${\cal O}(G^{2})$. However, imposing those conditions can lead to a consistent computation for both PM orders, which could manifest some generic features at different PM orders.
@@
$\partial_{r}U_{2}=$ $\frac{1}{2}{h^{(f)}_{2}}_{r_{f}r_{f}}-\frac{1}{2}(\partial_{r}U_{1})^{2}-\partial_{r}U_{1}\partial_{r}R_{1}+\frac{1}{2}r^{2}\bar{\gamma}_{AB}\partial_{r}X_{1}^{A}\partial_{r}X_{1}^{B}$
$+{h^{(f)}_{1}}_{r_{f}u_{f}}\partial_{r}U_{1}+{h^{(f)}_{1}}_{r_{f}r_{f}}\partial_{r}R_{1}+{h^{(f)}_{1}}_{r_{f}A}\partial_{r}X^{A}_{1}+\frac{1}{2}U_{1}\partial_{u}{h^{(f)}_{1}}_{r_{f}r_{f}}$
$+\frac{1}{2}R_{1}\partial_{r}{h^{(f)}_{1}}_{r_{f}r_{f}}+\frac{1}{2}X^{A}_{1}D_{A}{h^{(f)}_{1}}_{r_{f}r_{f}},$ (35)
$\partial_{r}R_{2}=$ ${h^{(f)}_{2}}_{u_{f}r_{f}}-\partial_{r}U_{1}\partial_{u}U_{1}-\partial_{r}U_{1}\partial_{u}R_{1}-\partial_{r}R_{1}\partial_{u}U_{1}+r^{2}\bar{\gamma}_{AB}\partial_{r}X_{1}^{A}\partial_{u}X_{1}^{B}$
$+{h^{(f)}_{1}}_{r_{f}u_{f}}\partial_{u}U_{1}+{h^{(f)}_{1}}_{r_{f}r_{f}}\partial_{u}R_{1}+{h^{(f)}_{1}}_{r_{f}A}\partial_{u}X_{1}^{A}+{h^{(f)}_{1}}_{u_{f}u_{f}}\partial_{r}U_{1}$
$+{h^{(f)}_{1}}_{u_{f}r_{f}}\partial_{r}R_{1}+{h^{(f)}_{1}}_{u_{f}A}\partial_{r}X_{1}^{A}+U_{1}\partial_{u}{h^{(f)}_{1}}_{u_{f}r_{f}}+R_{1}\partial_{r}{h^{(f)}_{1}}_{u_{f}r_{f}}$
$+X_{1}^{A}D_{A}{h^{(f)}_{1}}_{u_{f}r_{f}}-\partial_{r}U_{2}-\partial_{u}U_{2},$ (36)
@@
$+U_{1}\partial_{u}{h^{(f)}_{1}}_{{r_{f}B}}+R_{1}\partial_{r}{h^{(f)}_{1}}_{{r_{f}B}}+X_{1}^{C}D_{C}{h^{(f)}_{1}}_{{r_{f}B}}+{h^{(f)}_{1}}_{Bu_{f}}\partial_{r}U_{1}$
$+{h^{(f)}_{1}}_{Br_{f}}\partial_{r}R_{1}+{h^{(f)}_{1}}_{BC}\partial_{r}X_{1}^{C}-\partial_{r}U_{1}D_{B}U_{1}-\partial_{r}U_{1}D_{B}R_{1}-\partial_{r}R_{1}D_{B}U_{1}$
$-D_{B}U_{2}+2rR_{1}\partial_{r}X_{1B}\bigg)-\bar{\gamma}^{AB}\partial_{r}X_{1C}\partial_{B}X_{1}^{C}+X_{1}^{C}\partial_{C}\bar{\gamma}^{AB}\partial_{r}X_{1B}.$ (37)
@@
$+\frac{1}{2r^{2}}\bigg[{{h_{21}^{(f)}}_{r_{f}}}^{A}+{h_{11}^{(f)}}_{u_{f}r_{f}}D^{A}\beta_{1}-{{h_{1m}^{(f)}}_{B}}^{A}\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)$
$\qquad-2R_{10}\left({{h_{10}^{(f)}}_{r_{f}}}^{A}-D^{A}\beta_{1}\right)-D^{A}U_{21}+\beta_{1}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{A}$
$\qquad+\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)D_{B}{{h_{10}^{(f)}}_{r_{f}}}^{A}+{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{A}\left({{h_{10}^{(f)}}_{r_{f}B}}-D_{B}\beta_{1}\right)$
@@
$\qquad+\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{A}\left({{h_{11}^{(f)}}_{r_{f}B}}+\frac{1}{2}D_{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)+{{h_{11}^{(f)}}_{r_{f}B}}D^{A}\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)$
$\qquad+\frac{1}{2}\left({{h_{11}^{(f)}}_{r_{f}}}^{B}+\frac{1}{2}D^{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)D_{B}{{h_{10}^{(f)}}_{r_{f}}}^{A}+\left({{h_{10}^{(f)}}_{r_{f}}}^{B}-D^{B}\beta_{1}\right)D_{B}{{h_{11}^{(f)}}_{r_{f}}}^{A}$
@@
$\qquad{C_{2}}_{AB}={h_{2m}^{(f)}}_{AB}+2D_{(A}{h_{20}^{(f)}}_{B)r_{f}}-D_{A}D_{B}{h_{21}^{(f)}}_{r_{f}r_{f}}-2D_{A}D_{B}\beta_{2}+\bar{\gamma}_{AB}D^{2}\beta_{2}$
$\qquad\qquad\qquad-\frac{1}{2}\bar{\gamma}_{AB}\bar{\gamma}^{CD}{h_{2m}^{(f)}}_{CD}-\bar{\gamma}_{AB}D^{C}{h_{20}^{(f)}}_{r_{f}C}+\frac{1}{2}\bar{\gamma}_{AB}D^{2}{h_{21}^{(f)}}_{r_{f}r_{f}},$ (47)
@@
$+\frac{1}{3}D^{B}\beta_{1}\partial_{u}\left({h_{10}^{(f)}}_{AB}+D_{(A}{h_{10}^{(f)}}_{B)r_{f}}+\frac{1}{2}D_{A}D_{B}{h_{12}^{(f)}}_{r_{f}r_{f}}\right)$
$+D^{B}\beta_{1}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)+\frac{1}{2}D^{B}\beta_{1}D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)$
$+\frac{1}{2}D^{B}\beta_{1}D^{C}\left(D_{A}{h_{1m}^{(f)}}_{BC}-D_{B}{h_{1m}^{(f)}}_{AC}\right)$ (49)
@@
$-\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{C}\left(D_{A}{h_{1m}^{(f)}}_{BC}-D_{B}{h_{1m}^{(f)}}_{AC}\right)$
$-\frac{1}{2}{{h_{10}^{(f)}}_{r_{f}}}^{B}D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)$
$-\frac{1}{3}{{h_{10}^{(f)}}_{r_{f}}}^{B}\left(\partial_{u}{h_{10}^{(f)}}_{AB}+\partial_{u}D_{(A}{h_{10}^{(f)}}_{B)r_{f}}\right)-{{h_{10}^{(f)}}_{r_{f}}}^{B}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)$
$+\frac{1}{6}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{B}\left({h_{1m}^{(f)}}_{AB}+D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right).$
@@
$\begin{split}&D^{2}\left(D_{A}{h_{10}^{(f)}}_{r_{f}B}-D_{B}{h_{10}^{(f)}}_{r_{f}A}\right)\\ =&2D^{C}\left(D_{A}D_{(C}{h_{10}^{(f)}}_{B)r_{f}}-D_{B}D_{(C}{h_{10}^{(f)}}_{A)r_{f}}\right)+2\left(D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right).\end{split}$ (50)
@@
${N_{2}}_{A}\equiv$ ${h_{21}}_{uA}$
$=$ ${h_{21}^{(f)}}_{u_{f}A}+D_{A}{h_{22}^{(f)}}_{u_{f}r_{f}}+\frac{1}{3}\partial_{u}{h_{22}^{(f)}}_{r_{f}A}+\frac{1}{3}\partial_{u}D_{A}{h_{23}^{(f)}}_{r_{f}r_{f}}$
@@
$+\frac{1}{3}(D^{B}\beta_{1}-{{h_{10}^{(f)}}_{r_{f}}}^{B})\partial_{u}\left({h_{10}^{(f)}}_{AB}+D_{(A}{h_{10}^{(f)}}_{B)r_{f}}\right)$ (52)
@@
$+\frac{1}{6}\partial_{u}{{h_{11}^{(f)}}_{r_{f}}}^{B}\left({h_{1m}^{(f)}}_{AB}+D_{B}{h_{10}^{(f)}}_{r_{f}A}-D_{A}{h_{10}^{(f)}}_{r_{f}B}\right),$